OpenStudy (anonymous):

The hypotenuse of a right triangle is 25 cm long. One leg of the same triangle is 7 cm long. What is the length of the other leg? Explain how you found your answer.

OpenStudy (angela210793):

x^2=25^2-7^2=625-49=576 x=sqr576=24

OpenStudy (anonymous):

pythagoras theorem

OpenStudy (angela210793):

Pythagorean theorem

OpenStudy (anonymous):

hyp^2=base^2+perp^2